OpenStudy (anonymous):

Christina Hercher borrowed $50,000 on a 90 day, eight percent note. Christina paid $3,000 toward the note on day 40. On day 60 she paid an additional $4,000. Using the U.S. Rule, Christina's adjusted balance after the first payment is?

OpenStudy (anonymous):

What is the US Rule? I don't know if a=pe^rt applies here, but if so: 50000*e^(1.08*40/365)=56282.25 56282.25-3000=53282.25